2) Solve with UC superposition. y - 8y + 7y = e2x + 2x
Answer #1

АРЬ) Solve with UC Super position: y - syltry taa y- 84 + 7y Soln taa The Auxilary equation of given DE m2-8 M +7=0 m2-7m--53 = 1 } 61- 794 = 2 Cysg 7 0 70-804 cs 8 Cy 16 8 X 2 7x7 7 49 16 ملی 49 Now we put values of socy and es in eqh ( 22 P. I =